the 2. A rigid wire 40 oo cm current of 20.00 amps to the right. 15.00 cm, magnetic long a of y E field strength and orientation at a point wire is not infinitely long perpendicularly from the right end the wire. (this L 40 cm
Answer #1

given

l = 40 cm = 0.4 m

I = 20 A

y = 15 cm = 0.15 m


at point P, B = ?


we know,

B = (mue*I/(4*pi*y))*(cos(theta)) (here theta is the angle mae by the wire with +x axis)

= (mue*I/(4*pi*y))*(L/sqrt(L^2+y^2))

= (4*pi*10^-7*20/(4*pi*0.15))*(0.4/sqrt(0.4^2 + 0.15^2))

= 1.25*10^-5 T (out of the page)